a balloon was inflated to a volume of 5.0 L at a temperature of 2.5 degrees Celsius. It drifted over the mouth of an active volcano and expanded to 14.4L. What temperature did the air in the balloon reach when it entered the volcano?
Answer #1

volume V1 = 5.0 L

temperature T1 = 2.5 oC = 275.65 K

volume V2 = 14.4 L

temperature T2 = ??

V1 / T1 = V2 / T2

5.0 / 275.65 = 14.4 / T2

T2 = 794 K

temperature = 520.7 oC